OpenStudy (gabby1214):

solve using multiplication or division 4.4=-7+x x=?(Simplify answer, integer or decimal)

OpenStudy (benlindquist):

4.4 + -7 right

pooja195 (pooja195):

Flip the equation. x−7=4.4 Add 7 to both sides :) x=4.4+7 4.4+7=?

OpenStudy (kfireangel):

subract -7 from both sides, then x is alone and you have your answer

OpenStudy (gabby1214):

11.4?